Could you please solve 4(7x+10)=3(9x-5)+20? I really need someone to go through each step for me. I can show you what I tried. 4(7x+10)=3(9x-5)+20
28x+40=27x-15+20

Answer by stanbon(48545)

4(7x+10)=3(9x-5)+20
= 28x+40=27x-15+20
x=-35
Cheers,
Stan H.